OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of ?

OpenStudy (anonymous):

and i got 5y = 3x - 19

OpenStudy (anonymous):

but my choice are :

OpenStudy (anonymous):

5x – 3y = 19 5x – 3y = –19 3x – 5y = 19 3x – 5y = –19

OpenStudy (anonymous):

helppppppppppppppppp..

OpenStudy (anonymous):

what's the slope?

OpenStudy (anonymous):

3/5